# RoMedQA: The First Benchmark for Romanian Medical Question Answering

### v1

@@ 5.2 Preliminary Setup and Experiments
Table 6: Results of Phi-4-mini-instruct on RoMedQA, using only the first 2,048-token chunk of the epicrisis vs. all 2,048-token chunks. The former approach significantly outperforms the latter one across all metrics, suggesting that only analyzing the first part of the epicrisis reduces input clutter and prevents overfitting. The best results are highlighted in bold green.
@@ 5.3 Results
In Table 5, we present the results obtained by the chosen LLMs vs. two baselines on RoMedQA. We evaluate all four LLMs under both zero-shot prompting and supervised fine-tuning setups, reporting performance metrics on both validation and test sets.
@@
Use of multiple evaluation metrics. Regarding the employed evaluation metrics, we observe that EM scores exceed F1 scores by small margins across most configurations, suggesting that the models tend to either match answers exactly or miss them entirely, with fewer partial matches. Moreover, METEOR scores are consistently lower than other metrics, likely reflecting the sensitivity of METEOR to synonyms and word variations in Romanian medical terminology, but also its inherent bias against short answers. Nevertheless, for a holistic evaluation of QA models on RoMedQA, we promote the use of all four evaluation metrics.
Performance per question type. In Figure 3, we present the results of our best performing model, for each of the three question categories. The F1 and EM metrics suggest that binary questions are easier to answer, while reasoning questions are harder. This observation is consistent with our natural expectation. However, the BLEU score seems to promote reasoning questions, while METEOR indicates that the model is more capable on extractive questions. Overall, the analysis shows that some metrics are more appropriate for certain question categories. This represents another strong argument in favor of using a wide variety of metrics to accurately benchmark models on RoMedQA.
@@ 5.4 Qualitative Analysis
To better understand the challenges posed by RoMedQA and the capabilities of the best-performing model, we conduct a qualitative analysis of a random selection of QA pairs with correct and wrong answers predicted by Phi-4-mini-instruct. In Table 7, we showcase 10 examples, 5 that are correctly answered and 5 that are incorrectly answered, respectively.
@@ 6 Conclusion
In this paper, we introduced RoMedQA, the first benchmark for Romanian medical question answering. The new benchmark is the result of a meticulous data collection and annotation process carried out by experts in the medical domain, who constructed a dataset of 102,646 QA pairs. The QA samples are grounded in a collection of 1,011 medical case summaries of cancer patients. We conducted comprehensive experiments with several LLM families to assess the impact of various aspects, namely prompt structure and length, language vs. domain vs. general pretraining, and zero-shot prompting vs. LoRA-based fine-tuning. Our empirical results suggested some interesting findings. First of all, we observed that long prompts are not always helpful, degrading performance even if epicrises are typically long documents. Second of all, we found that generic pretraining can lead to better results, suggesting that prior language or domain adaption can actually produce LLMs with reduced generalization capabilities. Third of all, we determined that fine-tuning on our specific task significantly outperforms zero-shot prompting. Yet, even the best model was only able to correctly answer about $70\%$ of the test questions, indicating that RoMedQA is a challenging benchmark.

### v2

@@ 5.2 Preliminary Setup and Experiments
Table 6: Results of Phi-4-mini-instruct on MedQARo, using only the first 2,048-token chunk of the epicrisis vs. all 2,048-token chunks. The former approach significantly outperforms the latter one across all metrics, suggesting that only analyzing the first part of the epicrisis reduces input clutter and prevents overfitting. The best results are highlighted in bold green.
@@ 5.3 Results
In Table 5, we present the results obtained by the chosen LLMs vs. two baselines on MedQARo. We evaluate all four LLMs under both zero-shot prompting and supervised fine-tuning setups, reporting performance metrics on both validation and test sets.
@@
Use of multiple evaluation metrics. Regarding the employed evaluation metrics, we observe that EM scores exceed F1 scores by small margins across most configurations, suggesting that the models tend to either match answers exactly or miss them entirely, with fewer partial matches. Moreover, METEOR scores are consistently lower than other metrics, likely reflecting the sensitivity of METEOR to synonyms and word variations in Romanian medical terminology, but also its inherent bias against short answers. Nevertheless, for a holistic evaluation of QA models on MedQARo, we promote the use of all four evaluation metrics.
Performance per question type. In Figure 3, we present the results of our best performing model, for each of the three question categories. The F1 and EM metrics suggest that binary questions are easier to answer, while reasoning questions are harder. This observation is consistent with our natural expectation. However, the BLEU score seems to promote reasoning questions, while METEOR indicates that the model is more capable on extractive questions. Overall, the analysis shows that some metrics are more appropriate for certain question categories. This represents another strong argument in favor of using a wide variety of metrics to accurately benchmark models on MedQARo.
@@ 5.4 Qualitative Analysis
To better understand the challenges posed by MedQARo and the capabilities of the best-performing model, we conduct a qualitative analysis of a random selection of QA pairs with correct and wrong answers predicted by Phi-4-mini-instruct. In Table 7, we showcase 10 examples, 5 that are correctly answered and 5 that are incorrectly answered, respectively.
@@ 6 Conclusion
In this paper, we introduced MedQARo, the first large-scale benchmark for Romanian medical question answering. The new benchmark is the result of a meticulous data collection and annotation process carried out by experts in the medical domain, who constructed a dataset of 102,646 QA pairs. The QA samples are grounded in a collection of 1,011 medical case summaries of cancer patients. We conducted comprehensive experiments with several LLM families to assess the impact of various aspects, namely prompt structure and length, language vs. domain vs. general pretraining, and zero-shot prompting vs. LoRA-based fine-tuning. Our empirical results suggested some interesting findings. First of all, we observed that long prompts are not always helpful, degrading performance even if epicrises are typically long documents. Second of all, we found that generic pretraining can lead to better results, suggesting that prior language or domain adaption can actually produce LLMs with reduced generalization capabilities. Third of all, we determined that fine-tuning on our specific task significantly outperforms zero-shot prompting. Yet, even the best model was only able to correctly answer about $70\%$ of the test questions, indicating that MedQARo is a challenging benchmark.